OpenStudy (anonymous):

t^2dy/dt+2ty-y^5=0,t>0 solve the bernoulli equation by using a substitution v=y^1-n

OpenStudy (anonymous):

can any one help me plz

OpenStudy (anonymous):

ok, first of all, you need your first fan

OpenStudy (anonymous):

above is the question

OpenStudy (anonymous):

hmm... let me think

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nah, sorry, TOO hard for me, cant figure it out, I am not good with those derivatives and stuff, sorry....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you want to derive it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the question:\[dy/dt (t^2 +2ty - y^5)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

using v = y - n?

OpenStudy (anonymous):

i need to solve it by bernoulli equation substitution method

OpenStudy (anonymous):

v=y^1-n

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can take y = v + n

OpenStudy (anonymous):

substitute Y in the equation and you'll get :\[t^2 + 2t(v+n) - (v+n)^5\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you want to derive with respect to t?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you'll treat (v+n) in the equation as constants

OpenStudy (anonymous):

and so, I think you'll get : = 2t + 2(v+n) -(v +n) = 2t +2v + 2n -v -n = 2t +v +n I guess :)

OpenStudy (anonymous):

since you're deriving with respect to t, then you'll ignore the others and treat them as constants. please correct me if I'm wrong

myininaya (myininaya):

y and v are functions of t though. Also, yummy the site andrius gave you it awesome and easy to read

OpenStudy (anonymous):

lol yeah, google it :)